# AttentionPredictor: Temporal Pattern Matters for Efficient LLM Inference

### v2

@@ 4.2 AttentionPredictor: A Spatiotemporal Predictor
Prediction formulation. We formulate the attention history $A_{H}\in\mathbb{R}^{t\times t}$ as a spatiotemporal sequence. The first dimension of $A_{H}$ corresponds to the time series over the decoding steps, while the second dimension represents a sparse series over different keys. We then train a model to predict the attention for step $t$ as $\hat{A}_{t+1}=F(A_{H})$, where $F(\cdot)$ denotes the model function. For efficiency, we limit the time steps of $A_{H}$ using a hyperparameter $H$, so that the input to the predictor is $A_{H}\in\mathbb{R}^{H\times t}$.
Model design. To capture spatiotemporal features, we use a convolutional neural network (CNN) composed of two 2D convolution layers followed by a 1D convolution layer. The 2D convolutions capture spatiotemporal features at multiple scales, while the 1D convolution focuses on the time dimension, extracting temporal patterns across time steps. By replacing the fully connected layer with a 1D convolution kernel, the model adapts to the increasing spatial dimension, without data segmentation or training multiple models. Compared to an auto-regressive LSTM (Graves & Graves, 2012), the CNN is more lightweight and offers faster predictions, maintaining a prediction time shorter than the single-token inference latency. Additionally, when compared to an MLP (Rumelhart et al., 1986) on time-series dimension, the CNN is more effective at capturing spatial features, which improves prediction accuracy.
Training strategy. Our model is both data-efficient and generalizable. We train the model only on a small subset of attention data, specifically approximately 3% extracted from the dataset. The model performance on the entire dataset shows our model effectively captures the patterns (see Section 5.3). Additionally, due to the temporal characteristics of attention inherent in the LLM, a single model can generalize well across various datasets. For example, our model trained on LongBench also performs well on the GSM8K dataset, highlighting the generalization capability of AttentionPredictor.
Block-wise attention compression. To speed up prediction, we apply attention compression before computation. By taking advantage of the attention’s locality, AttentionPredictor predict attention and identify critical tokens in blocks. Inspired by Tang et al. (2024), we use the maximum attention value in each block as its representative. Specifically, max-pooling is applied on $A$ with a kernel size equal to the block size $b$, as $A_{t}^{comp}=Maxpooling(A_{t},b)$, reducing prediction computation to roughly $\frac{1}{b}$.
Distribution error calibration. Due to the sparsity of attention computation, the distribution of attention history $A_{H}$ used for prediction may deviate from the distribution of dense attention. This deviation tends to accumulate over decoding, particularly as the output length increases. To mitigate this issue and enhance prediction accuracy, we introduce a distribution error calibration technique to correct these deviations. Specifically, we calculate and store the full attention score every $M$ steps, effectively balancing accuracy with computational efficiency.
Overall process. As shown in Figure 3 and Algorithm 1, AttentionPredictor prepares an attention history queue in the prefilling stage, and predicts attention during the decoding stage. First, the $A_{t}$ from the LLM is compressed to $A_{t}^{comp}$ using block-wise attention compression. Next, $A_{H}$ is updated with $A_{t}^{comp}$. The next step attention $\hat{A}_{t+1}$ is then predicted with the pretrained model. From $\hat{A}_{t+1}$, the top-K positions are selected with a budget of $B/b$, since $\hat{A}_{t+1}$ is in compressed form. Finally, the indices are expanded with $b$ to obtain the final critical token positions $\mathbf{p}$.
@@
Input: Attention scores $A_{t}$, Attention history $A_{H}$, Block size $b$, KV budget $B$
Output: Critical KV token positions $\mathbf{p}$
1:  Pad $A_{t}$ to the nearest multiple of $b$ with zero
2:  $A_{t}^{comp}\leftarrow\text{MaxPooling}(A_{t},b)$
3:  $A_{H}\leftarrow\text{Update}(A_{h},A_{t}^{comp})$
4:  $\hat{A}_{t+1}\leftarrow\text{Prediction model}(A_{H})$
[…]
6:  $\mathbf{p}\leftarrow\text{Expand}(\text{Positions},b)$ Return positions $\mathbf{p}$
@@ 4.3 KV Cache Cross-token Prefetching
To address the increased memory cost of longer contexts, current LLM systems offload the KV cache to the CPU, but I/O transfer latency becomes the new significant bottleneck in inference. KV cache prefetching offers a solution by asynchronously loading important cache portions in advance, hiding retrieval time. We introduce the cross-token KV cache prefetching framework, which differs from the cross-layer method in Infinigen (Lee et al., 2024) by leveraging longer transfer times and enhancing data integration. Specifically, our implementation involves a prefetching process for each layer. As illustrated in Figure 3, during the prefill phase, the computed KV cache is completely offloaded to the CPU without compression. Then, AttentionPredictor forecasts the critical token indices $\mathbf{p}$ for the next step. The framework then prefetches the KV cache with $\mathbf{p}$ for the next step onto the GPU. Concurrently, the GPU processes inference for other layers, so the maximum time available for prediction and cache loading corresponds to the inference time per token. Subsequently, the GPU utilizes the query for the next step along with the prefetched partial KV cache to calculate the sparse attention. The attention history is then updated with the newly computed attention scores. The timeline of cross-token prefetching can be seen in Figure 4.
@@ 5.1 Settings
Tasks. We use the LongBench (Bai et al., 2024) dataset for long context evaluation. LongBench is a widely used benchmark for long-context LLM inference. It consists of 16 datasets covering 6 tasks, including single- and multi-document QA (SQA and MQA), summarization, few-shot learning, synthetic tasks, and code completion. LongBench uses different metrics for each dataset and calculates an average score to measure overall performance. See details in App. C.2.
We employ the GSM8K (Cobbe et al., 2021) dataset, a mathematics dataset annotated with CoT (Chain-of-Thought) reasoning processes, for evaluating CoT tasks. To evaluate the performance of our method with varying prompt lengths, we adjust the few-shot number. The evaluation metric is the accuracy of the final computational results.
Baselines. We select four sparse attention approaches as our baselines. These include three cache eviction methods, StreamingLLM (Xiao et al., 2024), H2O (Zhang et al., 2023), SnapKV (Li et al., 2024b) and one cache retrieval method Quest (Tang et al., 2024). For a fair comparison, SnapKV’s window size is set to 64, consistent with our setting. Aside from this modification, we use the official implementations and default parameters of the methods for reproducing the experiments. More details in App. C.1.
We choose two widely used long-context models for our evaluation: LongChat-v1.5-7b-32k (Kwon et al., 2023a) with 32K context length and LLaMA-3.1-8B-Instruct (Meta AI, 2024) with 128K context length.
Implementation details. Predictor Preparation. AttentionPredictor predictor consists of three convolutional layers, with $3\times 3$ kernels in the first two layers and a $1\times 1$ kernel in the last layer. For each LLM, the AttentionPredictor models are trained on a small subset of the datasets, where five samples (around 3%) were selected from each task in the LongBench dataset. During training, only decoding-stage attention was utilized. The trained AttentionPredictor models are evaluated on all tasks, including LongBench and GSM8K. The impressive results on GSM8K highlight the generalization capabilities and effectiveness of our approach.
@@
Hyper-parameters. We set the history step H to 64, the block size b to 16 and the calibration step M to 5. Performance analysis of these hyperparameters is discussed in subsection 5.4. Follow Quest (Tang et al., 2024), we did not apply our method or any other algorithms to the first two layers of the LLM. Following the settings of H2O and StreamingLLM, We allocated the budget equally to the prefix and local tokens, assigning 64 tokens each. The remaining KV budget is allocated to intermediate tokens, determined by the prediction model. We conducted experiments on NVIDIA A800 (80GB) GPUs.
5.2 Attention Recovery Rate
Table 1: The attention recovery rate (%) across different cache sizes (higher is better).
Llama-3.1-8B
KV Budget Method QA Summary Math Average
H2O 79.97 79.36 88.23 82.52
Quest 80.85 72.37 78.14 77.12
512 AttentionPredictor 84.79 79.73 87.77 84.10
H2O 85.70 83.49 91.37 86.85
Quest 87.69 81.02 85.52 84.74
1024 AttentionPredictor 90.63 85.16 92.04 89.28
H2O 91.58 88.37 94.79 91.58
Quest 93.40 88.12 92.04 91.19
2048 AttentionPredictor 95.03 90.30 95.52 93.62
H2O 96.59 93.49 97.97 96.02
Quest 97.37 93.93 97.06 96.12
4096 AttentionPredictor 98.07 94.88 98.26 97.07
We first evaluate the attention recovery rate $R_{rec}$ of AttentionPredictor, defined as the proportion of attention scores over critical tokens identified by the method relative to the total scores as defined in Equation 4, which reflects the intermediate information loss caused by KV cache compression methods. On the three representative tasks—QA, summary, and mathematical reasoning—with different KV cache budgets, AttentionPredictor consistently achieves a higher average recovery rate compared to H2O(Zhang et al., 2023) and Quest (Tang et al., 2024) as shown in Table 1. This demonstrates that our method accurately identifies the positions of critical tokens, thereby minimizing information loss. Notably, with the KV budget of 512, AttentionPredictor shows a significant advantage over Quest, achieving a 7% higher average recovery rate, emphasizing our robustness under the extremely high (20$\times$) compression ratio. Furthermore, in the mathematical reasoning task, block-retrieval method Quest shows substantial recovery loss, suggesting potential limitations in this complex task. In contrast, our block-prediction method alleviates this issue, further underscoring its reliability.
5.3 Final Accuracy Evaluation
Results on LongBench. We evaluate our method on various long-context tasks in the LongBench benchmark, with the KV cache budgets ranging from 512 to 4096. We restrict H2O’s attention to the past 64 steps to ensure fairness and denote this variant as H2O+. As shown in Table 2, AttentionPredictor surpasses the performance of all SOTA KV cache eviction and retrieval methods across various KV budgets and LLMs. Notably, the average performance loss compared to the full cache is less than 0.5% across all cache budgets. This demonstrates the ability of our method to effectively model attention patterns and precisely predict the locations of critical tokens. Furthermore, under an extremely sparse budget setting of 4% (512/13K), our approach results in only a 0.44% decrease in performance compared to the full cache, indicating a 76% improvement over the state-of-the-art Quest method, which suffers a 1.84% reduction.
@@ 5.3 Final Accuracy Evaluation
Table 2: The evaluation results from the LongBench dataset across 512, 1024, 2048, and 4096 KV cache budgets.
@@
4096 AttentionPredictor 30.78 24.50 24.65 64.19 14.75 54.85 36.10 43.14 45.45 25.94 68.07 54.41 51.21 47.92
Results on CoT reasoning. As shown in Table 3, we evaluate our method on the mathematical reasoning dataset GSM8K with LLaMA-3.1. By adjusting the number of few-shot examples to simulate the long-context reasoning process with short and long CoTs, we generate input lengths of 4K, 8K, and 16K. Unlike long-context tasks, CoT mathematical reasoning tasks present distinct challenges. For example, the retrieval-based SOTA method, Quest, excels on long-context benchmarks but performs poorly on CoT tasks, showing a 16.91% accuracy drop at a sequence length of 16K. In contrast, AttentionPredictor achieves a significantly smaller accuracy loss of just 2.05%. Moreover, across all sequence lengths, our method outperforms baselines in most cases. Under a 25% cache budget, it even exceeds the accuracy of the full cache, achieving lossless performance. H2O+ also performs well in this task, likely due to half of the cache budget being on local tokens, which remain unaffected by variations of the few-shot. Additionally, H2O+ includes only 64 steps of historical information, limiting the influence of irrelevant attention scores.
Table 3: Evaluation results on the CoT dataset.
GSM8K+LLaMA-3.1
Method Budget Prompt length
4K 8K 16K Average
Full cache Full 56.79 55.27 54.13 55.40
StreamingLLM 1K 54.74 49.96 50.94 51.88
H2O+ 1K 57.16 52.01 52.16 53.78
Quest 1K 48.52 45.26 37.22 43.67
SnapKV 1K 53.45 48.67 49.20 50.44
AttentionPredictor 1K 57.16 53.75 52.08 54.33
@@ 5.4 Ablation Study
The effects of varying hyperparameters on the performance of AttentionPredictor with Longchat are studied. Three key hyperparameters, including history step, calibration step, and block size, are evaluated across six tasks from the LongBench dataset. Detailed task-specific results are available in App. C.5. Additionally, the prediction model implementation is also evaluated.
Block size $b$. As shown in 5(a), AttentionPredictor maintains superior average performance across block sizes from 8 to 64. While performance degradation occurs with increasing block size due to coarser token positioning, our method exhibits a milder decline compared to Quest’s block-granularity KV cache retrieval. The performance drop in Quest may be attributed to the loss of total attention score information when the min-max based attention score upper bound estimation of each block is used for retrieved, leading to inaccuracies in block identification. In contrast, our approach accurately captures attention patterns, thereby improving critical block identification and mitigating the drawbacks of block-wise retrieval.
Calibration step $M$. 5(b) shows the average performance of AttentionPredictor with calibration steps ranging from 1 to 20. Performance improves with shorter calibration intervals, indicating that the calibration scheme reduces cumulative errors caused by differences between sparse and original attention distributions. However, higher calibration frequencies increase computational costs, creating a trade-off between accuracy and efficiency.
History step $H$. As depicted in 5(c), we evaluate the average accuracy of AttentionPredictor over a range of history steps. Overall, the performance gap compared to the full cache remains consistently small. Notably, the middle value of $H$ maximizes performance by providing the necessary information for pattern recognition, while mitigating redundancy caused by the decaying self-correlation of attention scores.
(a) Block Size
(b) Calibration Step
(c) History Step
Figure 5: Hyper-parameter ablation on Longchat with LongBench.
Prediction model. We evaluate the performance of various prediction model implementations. MLP is applied to the sequence dimension to handle variations in attention length. For LSTM, attention is divided into 16 width blocks and predictions are independent for each block, referred to as LSTM-block in Table 4. Similarly, CNN-block utilizes a block-wise prediction approach, employing a fully connected layer for fixed block lengths. Finally, the CNN model, as the primary setting in our experiments, predicts all attention scores simultaneously. As reported in Table 4, CNN outperforms other models, achieving the highest attention recovery rate by effectively capturing attention patterns. MLP failed to account for neighboring token interactions, while LSTM-block and CNN-block were restricted to block-level information without global context. Notably, CNN also required the fewest parameters and was the most memory-efficient during inference.
5.5 Efficiency
Table 4: Attention recovery rate and parameter numbers of different model implementation.
LongBench+Longchat
Model type Total params Recovery rate
MLP 49K 92.88
LSTM-block 3.2M 91.75
CNN-block 21K 91.94
CNN 4.9K 93.65
We evaluate the final LLM inference acceleration of our method in Table 5, using the widely adopted cross-layer prefetching system with cache offloading as the baseline, which is implemented in Wolf et al. (2020). By reducing cache transfers, AttentionPredictor achieves a 1.2$\times$ speedup in per-token decoding with a 1K budget, demonstrating more efficient inference. The added token evaluation time is negligible, leading to a substantial overall speed improvement. Furthermore, we apply AttentionPredictor to our proposed cross-token prefetching framework, which enables asynchronous cache retrieval and I/O transferring between layers. This integration yields an additional 1.4$\times$ speedup in the decoding stage, further highlighting the effectiveness of our approach.
Table 5: The decode latency of per token of LLaMA-3.1-8B with 32K context length.
Method Decode latency(s) Speed
cross-layer pref. 0.364 1$\times$
cross-layer pref.+AttentionPredictor 0.295 1.2$\times$
cross-token pref.+AttentionPredictor 0.262 1.4$\times$

### v3

@@ 3.4 AttentionPredictor: A Spatiotemporal Predictor
Overall Process. As shown in Figure 3 and Algorithm 1, AttentionPredictor prepares an attention history sequence $\mathcal{A}_{H}$ in the prefilling stage, and predicts attention during the decoding stage. First, the $A_{t}$ from the LLM is compressed to $A_{t}^{comp}$ using block-wise attention compression. Next, $\mathcal{A}_{H}$ is updated with $A_{t}^{comp}$. The next step attention $\hat{A}_{t+1}$ is then predicted with the pretrained model $\mathcal{F}$. From $\hat{A}_{t+1}$, the top-K positions are selected with a budget of $B/b$, since $\hat{A}_{t+1}$ is in compressed form. Finally, the indices are expanded with $b$ to obtain the final critical token indices $S$.
Input: Attention scores $A_{t}$, Attention history series $\mathcal{A}_{H}$, Block size $b$, KV budget $B$
Output: Critical KV token indices set $S$
1: Pad $A_{t}$ to the nearest multiple of $b$ with zero
[…]
Return KV cache indices set $S$
@@
Model Design. To capture spatiotemporal features, we introduce a convolutional neural network (CNN) composed of two 2D convolution layers followed by a 1D convolution layer. The 2D convolutions capture spatiotemporal features at multiple scales, while the 1D convolution focuses on the time dimension, extracting temporal patterns across time steps. By replacing the fully connected layer with a 1D convolution kernel, the model adapts to the increasing spatial dimension, without data segmentation or training multiple models. Compared to an auto-regressive LSTM [43], the CNN is more lightweight and offers faster predictions, maintaining a prediction time shorter than the single-token inference latency. Additionally, when compared to a MLP [44] on time-series dimension, the CNN is more effective at capturing spatial features, which improves prediction accuracy.
Training Data and Strategy. Our model is both data-efficient and generalizable. We train the model only on a small subset of attention data, specifically approximately 3% extracted from the dataset. The model performance on the entire dataset shows our model effectively captures the patterns (see Section 4.2). Additionally, the temporal patterns of attention are inherent in LLMs, a single model can generalize well across various datasets. For example, our model trained on LongBench also performs well on the GSM8K dataset, highlighting the generalization capability of AttentionPredictor.
Block-wise Attention Compression. To speed up prediction, we apply attention compression before computation. By taking advantage of the attention’s locality, AttentionPredictor predict attention and identify critical tokens in blocks. Inspired by Quest [31], we use the maximum attention value in each block as its representative. Specifically, max-pooling is applied on $A_{t}$ with a kernel size equal to the block size $b$, as $A_{t}^{comp}=\text{Maxpooling}(A_{t},b)$, reducing prediction computation to roughly $1/b$.
Distribution Error Calibration. Due to the sparsity of attention computation, the distribution of attention history $\mathcal{A}_{H}$ used for prediction may deviate from the distribution of dense attention. This deviation tends to accumulate over decoding, particularly as the output length increases. To mitigate this issue and enhance prediction accuracy, we introduce a distribution error calibration technique to correct these deviations. Specifically, we store the full attention score every $M$ steps, effectively balancing accuracy with computational efficiency. Note that the full attention is only used to update the history. During LLM decoding, we still use the sparse KV cache, strictly adhere to the budget. Additionally, we predict and update the position $S$ every several steps for efficiency.
@@ 3.5 KV Cache Cross-token Prefetching
To address the increased memory cost of longer contexts, current LLM systems offload the KV cache to the CPU, but I/O transfer latency becomes a significant bottleneck in inference. KV cache prefetching offers a solution by asynchronously loading the sparse cache in advance. We introduce the cross-token KV cache prefetching framework, which differs from the cross-layer method [19] by leveraging longer transfer time budgets and enhancing data integration. They are helpful for improving prediction accuracy with larger models and improving transfering efficiency. Specifically, our implementation involves a prefetching process for each layer. As illustrated in Figure 3, during the decoding phase, AttentionPredictor forecasts the critical token indices $S$ for the next step. The framework then prefetches the sparse KV cache with $S$ for the next step from the full cache to the GPU. At the same time, LLM inference of other layers is also ongoing. Subsequently, the prefetched sparse KV cache is used to calculate the sparse attention of the next token. The timeline of cross-token prefetching can be seen in Figure 10 in Appendix B.1.
@@ 4.1 Settings
Tasks. 1) We use the LongBench [45], InfiniteBench [46] and RULER QA [47] dataset for long context evaluation. LongBench is a widely used benchmark for long-context LLM inference. It consists of 16 datasets covering 6 tasks, including single- and multi-document QA (SQA and MQA), summarization, few-shot learning, synthetic tasks, and code completion. The metrics for each task are in Appendix B.3. InfiniteBench has up to extreme 124K average context length. 2) We employ the AIME [48] dataset, a tough mathematics dataset, for evaluating the performance of our method with the long-reasoning output scenario. 3) We use GSM8K math dataset to evaluate the long n-shot CoT task. 4) We use MMLU [49] and GPQA [50] dataset to evaluate the multi-choice task. 5) We use the Needle In A Haystack [51] experiment to evaluate the in-context retrieval capabilities.
Baselines and LLMs. We select four sparse attention approaches as our baselines. These include four cache eviction methods, StreamingLLM [12] and its advanced method Cascading [52], accumulative approaches H2O [11] and SnapKV [13], and two cache retrieval method Quest [31] and learning-based SeerAttention [14]. We use the official implementations of all baseline experiments. A detailed description of each baseline is provided in Appendix B.2. We choose two widely used long-context models for our evaluation: LongChat-v1.5-7b-32k [53] with 32K context length and LLaMA-3.1-8B-Instruct [54] with 128K context length.
@@
Implementation details. We set the history step H to 64, the block size b to 16, and the calibration step M to 5. Performance analysis of these hyperparameters is discussed in Section 4.6. Follow Quest [31], we did not apply our method or any other algorithms to the first two layers of the LLM. Following the settings of H2O and StreamingLLM, We allocated the budget equally to the prefix and local tokens, assigning 64 tokens each. The remaining KV budget is allocated to intermediate tokens, determined by the prediction model. We conducted experiments on NVIDIA A800 (80GB) GPUs.
4.2 Main Results
@@
Table 1: The evaluation results on the LongBench dataset across 1K, 2K, and 4K KV cache budgets.
@@
4096 AttentionPredictor 30.78 24.50 24.65 64.19 14.75 54.85 36.10 43.87 44.57 25.69 69.07 53.38 53.44 48.17
Results on LongBench. We evaluate our method on various long-context tasks in the LongBench benchmark, with the KV cache budgets ranging from 1024 to 4096. We restrict H2O’s attention to the past 64 steps to ensure fairness and denote this variant as H2O+. As shown in Table 1, AttentionPredictor surpasses the performance of all SOTA KV cache eviction and retrieval methods across various KV budgets and LLMs. Notably, the average performance loss compared to the full cache is less than 0.5% across all cache budgets. This demonstrates the ability of our method to effectively model attention patterns and precisely predict the locations of critical tokens. Furthermore, in an extremely sparse budget setting of 8% (1K/13K), our approach achieves full cache performance, while all other methods suffer at least a 1.25% reduction. Additional results of SeerAttention, long-context benchmark InfiniteBench and RULER QA are in Appendix C.2, C.3, and C.4.
Figure 4: Evaluation results on the long-output reasoning task AIME2024 with QwQ-32B.
Figure 5: Per-token decode latency of LLaMA-3.1-8B under varying tokens, comparing ours with FlashAttention2 (full cache) and H2O.
Results on CoT reasoning. Reasoning significantly enhances the problem-solving capabilities of LLMs. For instance, QwQ-32B achieves an accuracy of 76.7% on the challenging AIME benchmark with the aid of detailed reasoning. However, generating complex answers leads to output length of up to 31.5K tokens. This sharply contrasts with typical long-context scenarios, which involve long inputs but short outputs. Such a shift introduces distinct challenges for KV cache compression during decoding. Under this demanding setting, we evaluate AttentionPredictor against two decoding-phase acceleration baselines, H2O and Quest. As shown in Figure 5, AttentionPredictor consistently outperforms both baselines. It achieves comparable accuracy of 76.7% with a KV budget ratio of 0.13 (2K/15K average), whereas Quest drops sharply to 43.3%. Moreover, our approach attains this accuracy with an average output of only 15K tokens, offering significantly higher efficiency compared to H2O (21K) and Quest (20K). Additional results of GSM8K, multi-choice benchmarks MMLU and GPQA are in Appendix C.1, C.5, and C.6.
4.3 Needle In A HayStack
We conduct the "Fact Retrieval Across Context Lengths" (Needle In A Haystack) experiment. Our approach achieved a perfect 100% retrieval accuracy with a 64K-token context and a 1/64 cache compression ratio, surpassing all existing methods. This outcome underscores the effectiveness of our method in preserving critical information necessary for accurate retrieval.
Figure 6: Results of the Fact Retrieval Across Context Lengths (“Needle In A HayStack”) test in LLaMA-3.1-8B-Instruct with 64k context size in 1024 KV cache size.
4.4 Efficiency
Decode latency. We evaluate decoding latency in a cache offloading setup [55], essential for long contexts where the KV cache exceeds GPU memory. In this scenario, the standard approach (FlashAttention2) creates a bottleneck by prefetching the entire KV cache from CPU to GPU, which our method optimizes. Further details on the setup are in Appendix B.4. As shown in Figure 5, our method consistently outperforms both FlashAttention2 and H2O across all context lengths. Notably, at 32K tokens, we achieve a 5.6× acceleration compared to FlashAttention2, demonstrating the dramatic acceleration enabled by loading only a small fraction of the cache. Moreover, we attain a 1.6× speedup over the calculation-efficient H2O, showing the effectiveness of our cross-token prefetch system.
Table 2: Per-layer overhead breakdown (in ms) of our cache prefetching system, averaged across all layers.
Context Length 4K 8K 16K 32K
Predict Attention 0.7 1.2 2.3 4.5
Transfer Cache 2.3 3.9 7.6 13.2
Wait for Main LLM 44.4 43.4 39.7 32.3
Total Per-Layer Overhead 47.3 48.6 49.6 50.0
Overload breakdown. Table 2 details the per-layer overhead of our prefetching system. A key result is that this total overhead remains remarkably stable, increasing only from 47.3ms to 50.0ms as context length grows from 4K to 32K. This stability is a core advantage of our asynchronous design. The growing latency from Predict Attention and Transfer Cache is effectively hidden by being absorbed into the Wait for Main LLM period, during which the main model is busy processing the current token’s inference. Due to parallel execution across layers, this per-layer overhead is equivalent to the final per-token latency.
4.5 Prediction Accuracy
[…]
Table 3: The attention prediction accuracy (%) across different cache sizes.
Llama-3.1-8B
KV Budget Method QA Summary Math Average
H2O 85.90 87.89 93.89 89.23
Quest 86.84 80.15 83.15 83.38
512 AttentionPredictor 91.07 88.30 93.40 90.93
H2O 89.79 89.92 95.40 91.70
Quest 91.87 87.26 89.29 89.47
1024 AttentionPredictor 94.95 91.72 96.10 94.25
H2O 93.82 92.61 97.09 94.51
Quest 95.69 92.35 94.27 94.10
2048 AttentionPredictor 97.36 94.63 97.84 96.61
H2O 97.37 95.65 98.75 97.26
Quest 98.16 96.10 97.83 97.36
4096 AttentionPredictor 98.86 97.07 99.04 98.33
Figure 7: The prediction accuracy (%) of different dataset from LongBench.
Generalization across datasets. We evaluate our model’s generalization by training exclusively on a single dataset and then evaluating prediction accuracy on all remaining, unseen tasks within LongBench. As shown in Figure 7, the model consistently achieves over 95% accuracy across every task, demonstrating strong cross-task transferability and generalization.
@@ 4.6 Ablation Study
The effects of hyperparameter block size on AttentionPredictor performance are studied. Other hyperparameters, predictor structure and training efficiency are shown in App. C.8, C.7, and D.1.
Figure 8: Evaluation of Block Size on Longchat with LongBench.
Block size $b$. Figure 8 shows that AttentionPredictor maintains superior average performance with block sizes 8-64. While larger block sizes degrade performance from coarser token positioning, AttentionPredictor declines more mildly than Quest’s block-wise KV cache retrieval. Quest’s performance drop may stem from losing total attention score information when its min-max based per-block attention score upper bound estimation is used for retrieval, causing inaccurate block identification. In contrast, AttentionPredictor accurately captures attention patterns, improving critical block identification and mitigating block-wise retrieval drawbacks.